OpenStudy (anonymous):

Solve for x: 2 over quantity x minus 2 plus 7 over quantity x squared minus 4 equals 5 over x

OpenStudy (anonymous):

@ngott3

OpenStudy (anonymous):

\[\frac{ 2 }{ x-2 } + \frac{ 7 }{ x^2 - 4 } = \frac{ 5 }{ x }\]

OpenStudy (anonymous):

Answer Choices: x = negative 4 over 3 and x = –5 x = negative 4 over 3 and x = 5 x = 4 over 3 and x = –5 x = 4 over 3 and x = 5

OpenStudy (anonymous):

ok for X^2-4 what can u expand that to?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):